OpenStudy (anonymous):

The perimeter of a rectangle is 68 feet. The length of the rectangle is 4 feet more than twice the width. Find the length and the width of the rectangle. Explain your reasoning.

OpenStudy (anonymous):

Perimeter = 68 or P = 68 P = 2L + 2W (twice the length plus twice the width) P = 2(4+2W) + 2W P = 8 + 6W 60 = 6W 10 = W then 68 = 2L + 2W 68 = 2L + 20 24 = L width = 10 length = 24